# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

### v2

@@ A.3 Information-Theoretic Lower Bound on Quantum Kolmogorov Complexity
Lemma A.3 (Information-Theoretic Lower Bound). For any $\varepsilon>0$, there exists a constant $c_{0}=O(1)$, which depends only on the relative universal machine $U_{R}$ and is independent of $d_{R}$, such that [24][23] [26]:
$K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$ (A.3)
@@ A.4 Entropy Identity
$D_{\max}(\rho||\sigma_{R})-D(\rho||\sigma_{R})=S(\rho)-H_{\mathrm{min}}(\rho)=\Delta_{\mathrm{spec}}(\rho)$ (A.4)
@@ A.5 Combining and Rearranging the Inequality Chain
$D_{\max}^{\varepsilon}(\rho||\sigma_{R})\geq D_{\max}(\rho||\sigma_{R})-\alpha(\varepsilon)$ (A.5)
Bounds of this type can be derived from the standard relationships between smoothed measures and the purified distance [50]. Second, by combining Lemma A.2 (Eq. (A.2)) and Lemma A.3 (Eq. (A.3)), we obtain the core inequality from physical cost to a smoothed information quantity. To ensure mathematical rigor, we uniformly use the form
$\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}):C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\cdot\log\Gamma_{R}+c_{a}\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+c_{b}\log(1/\varepsilon)+O(1)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$ (A.6)
Next, applying the aforementioned smoothing continuity bound (in the lower-bound direction) to the right-hand side and substituting the entropy identity proven in Proposition A.4 (Eq. (A.4)), we get:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\cdot\log\Gamma_{R}\geq D(\rho||\sigma_{R})+\Delta_{\mathrm{spec}}(\rho)-\alpha(\varepsilon)-c_{0}-c_{a}\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})-c_{b}\log(1/\varepsilon)-O(1)$ (A.7)
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{D(\rho||\sigma_{R})}{\log\Gamma_{R}}-\frac{c_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]+\frac{\Delta_{\mathrm{spec}}(\rho)}{\log\Gamma_{R}}$ (A.8)
@@ A.6.1 Key One-Shot Information-Theoretic Quantities
We first precisely define the core information-theoretic quantities on which the proof in this section relies. By convention: all logarithms $\log$ are base-2, errors are measured by the trace distance $(1/2)\cdot||.||_{1}$, and it is assumed by default that the support of all quantum states $\rho$ is contained within the subspace $H_{R}$ defined by the reference set RR[26].
@@
    $D_{H}^{\eta}(\rho||\sigma)=-\log\min_{0\leq T\leq I,\mathrm{Tr}[T\sigma]\leq\eta}\mathrm{Tr}[(I-T)\rho]$ (A.9)
@@
    $D_{\max}(\rho||\sigma)=\inf\{\lambda:\rho\leq 2^{\lambda}\cdot\sigma\}$ (A.10)
@@ A.6.2 Core Lemmas and Relations
$D_{H}^{\eta}(E(\rho)||E(\sigma))\leq D_{H}^{\eta}(\rho||\sigma)$ (A.12)
$D_{\max}^{\varepsilon}(E(\rho)||E(\sigma))\leq D_{\max}^{\varepsilon}(\rho||\sigma)$ (A.13)
@@
$D_{H}^{\eta}(\rho||\sigma)\leq D_{\max}^{\varepsilon}(\rho||\sigma)+\kappa_{1}\cdot\log(1/\eta)+\kappa_{2}\cdot\log(1/\varepsilon)$ (A.14)
@@ A.6.3 Upper Bound on Smoothed Max-Relative Entropy from Circuit Complexity
$D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq L\cdot\log\Gamma_{R}+a\cdot\log(\max\{1,L\})+b\cdot\log(1/\varepsilon)$ (A.15)
Proof. Following the prefix-free encoding idea from Section A.2, we construct the universal semi-density operator $M=\sum_{p}2^{-|p|}\rho_{p}$. If there exists a circuit described by a program $p^{*}$ that can $\varepsilon$-approximately prepare $\rho$ (with output $\rho_{p^{*}}$), then by the definition of $M$, we have the operator inequality $M\geq 2^{-|p^{*}|}\cdot\rho_{p^{*}}$, which directly leads to $D_{\max}(\rho_{p^{*}}||M)\leq|p^{*}|$. To change the comparison object from $M$ to $\sigma_{R}$, we introduce a mixed semi-density operator $M_{\alpha}=(1-\alpha)M+\alpha\sigma_{R}$, where $\alpha\in(0,1)$. Using the scaling comparison property of max-relative entropy, we get $D_{\max}(\rho_{p^{*}}||\sigma_{R})\leq D_{\max}(\rho_{p^{*}}||M)+\log(1/(\alpha(1-\alpha)))$. Choosing $\alpha=1/2$ makes the constant term $\log(1/(\alpha(1-\alpha)))=\log(4)=2$. Combining this with the information-theoretic upper bound on $|p^{*}|$, which is $L\cdot\log\Gamma_{R}+O(\log L)$, and the continuity of smoothed entropies, we obtain Eq. (A.15). This proof path, via the $\sigma$-padding technique, avoids any spectral constants with explicit dependence on $d_{R}$.
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\frac{1}{\log\Gamma_{R}}D_{H}^{\eta}(\rho||\sigma_{R})-\frac{c^{\prime}_{1}}{\log\Gamma_{R}}\left[\log(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\})+\log(1/\varepsilon)\right]-\frac{c^{\prime}_{2}}{\log\Gamma_{R}}\log(1/\eta)$ (A.16)
Proof. Let $L=C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$. We chain together Lemma A.6.2 and Lemma A.6.3. From Eq. (A.14) and Eq. (A.15), we have: $D_{H}^{\eta}(\rho||\sigma_{R})\leq\left(L\cdot\log\Gamma_{R}+a\cdot\log(\max\{1,L\})+b\cdot\log(1/\varepsilon)\right)+\kappa_{1}\cdot\log(1/\eta)+\kappa_{2}\cdot\log(1/\varepsilon)$ By algebraically rearranging the above expression to solve for a lower bound on $L$, and defining the universal constants $c^{\prime}_{1}=\max\{a,b+\kappa_{2}\}$ and $c^{\prime}_{2}=\kappa_{1}$, we obtain Eq. (A.16).
@@ A.6.5 Non-Bootstrap Closed-Form Lower Bound
Eq. (A.16) is a bootstrap-type inequality. For practical applications, we can provide a non-recursive, closed-form lower bound for it using the method derived in Appendix C.2. Let:
$A$ $=\frac{D_{H}^{\eta}\!\bigl(\rho\,\|\,\sigma_{R}\bigr)}{\log\Gamma_{R}},$ (A.17)
$B$ $=\frac{1}{\log\Gamma_{R}}\!\left(c^{\prime}_{1}\log\!\frac{1}{\varepsilon}\;+\;c^{\prime}_{2}\log\!\frac{1}{\eta}\right),$ (A.18)
$c$ $=\frac{c^{\prime}_{1}}{\log\Gamma_{R}}$ (A.19)
Then Eq. (A.16) can be written as $L\geq A-B-c\cdot\log(\max\{1,L\})$. Its smallest solution $L^{*}$ is given by the Lambert-W function $W_{0}$:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\geq\max\left\{1,\left(\frac{c}{\ln 2}\right)\cdot W_{0}\left(\left(\frac{\ln 2}{c}\right)\cdot\exp\left(\left(\frac{\ln 2}{c}\right)\cdot(A-B)\right)\right)\right\}$ (A.20)
@@ A.6.6 Framework Consistency and Independence Check
Accounting for Errors and Compilation: The $O(\log(1/\varepsilon))$ corrections arising from smoothing and gate synthesis have been uniformly incorporated into the logarithmic correction term of Eq. (A.16); the constant factor from compiling parallel to serial circuits does not affect the scaling behavior and only influences the specific numerical values of the $O(1)$ coefficients $a$ and $b$.
@@ B.1 Robustness of the Theory to Support Set Leakage
$\tilde{\rho}=\frac{\Pi_{R}\cdot\rho_{\mathrm{phys}}\cdot\Pi_{R}}{\mathrm{Tr}(\Pi_{R}\cdot\rho_{\mathrm{phys}})}$ (B.1)
All conclusions in the main theorem, Theorem 3.1, can be directly applied to this effective state $\tilde{\rho}$. The physical meaning of this scheme is that we focus our analysis on the part of the quantum state that was successfully prepared within the target structure, treating it as the true computational goal. (Note: The above construction requires $q=\mathrm{Tr}(\Pi_{R}\rho_{\mathrm{phys}})>0$. When $q=0$ (complete leakage), the smoothed reference state scheme (B.2) should be used instead.)
@@
$\tilde{\sigma}_{R}=(1-\delta)\cdot\sigma_{R}+\delta\cdot\frac{(I-\Pi_{R})}{(d-d_{R})}$ (B.2)
@@ B.2 Proof of the $d_{R}$-Independence of the Kolmogorov Lower Bound Constant $c_{0}$
Proposition B.3 ($d_{R}$-Independence of $c_{0}$). There exists a constant $c_{0}=O(1)$, whose magnitude depends only on the choice of the relative universal machine $U_{R}$, such that for any $\varepsilon>0$ and any $d_{R}$, it holds that $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$[51].
@@
    $D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq D_{\max}^{\varepsilon}(\rho||M)+\log(1/(\alpha(1-\alpha)))$ (B.4)
@@
    Combining and Parameter Selection: From Eq. (B.3) and Eq. (B.4), we have:
@@
    $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-2$ (B.5)
@@
(Note: We do not rely on any spectral comparison constants between $M$ and $\sigma_{R}$; using $\sigma$-padding (with $\alpha=1/2$) is sufficient to obtain $K_{\varepsilon}(\rho\!\mid\!R)\geq D_{\max}^{\varepsilon}(\rho\|\sigma_{R})-c_{0}$, where $c_{0}=O(1)$ and is independent of $d_{R}$, see Appendix A.3 and Section B.2.) [51]
@@ B.3 Comparison with and Positioning Relative to Other Complexity Measures
Table 1: Table B.1: A Multi-dimensional Comparison of RCC with Other Complexity Measures
Dimension Reference-Contingent Complexity (RCC) Nielsen’s Geometric Complexity Krylov Complexity Resource-Theoretic Relative Entropy Mana/T-count
Object of Measure State function (final state) Path function (dynamical process) Operator growth (dynamics) Resource content of a state Resource count for a specific gate set
Structural Fairness Explicitly modeled via reference set Depends on the choice of penalty factors Partial (depends on basis choice) Yes (depends on the definition of the free set) No (model is fixed)
Relation to Gate Set Explicitly dependent via $\Gamma_{R}$ Implicitly dependent via metric Typically independent of gate set Related to the set of free operations Strongly dependent on a specific gate set
Bound on Circuit Complexity Theorem-level lower bound Typically an upper bound or an approximation Indirectly related Not a direct bound Strict lower bound for a specific model
Operationality Yes (protocol in Sec. B.3.2) Indirect (requires path optimization) Yes (measurable via response functions) Partial (depends on free operations) Yes (can be counted directly by compiling)
Path Dependence No Yes Yes No No

### v3

@@ A.3 Information-Theoretic Lower Bound on Quantum Kolmogorov Complexity
Lemma A.3 (Information-Theoretic Lower Bound). For any $\varepsilon>0$, there exists a constant $c_{0}=O(1)$, which depends only on the relative universal machine $U_{R}$ and is independent of $d_{R}$, such that [berthiaume2001quantum][Gacs2001] [Tomamichel2012]:
$K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$ (33)
@@ A.4 Entropy Identity
$D_{\max}(\rho||\sigma_{R})-D(\rho||\sigma_{R})=S(\rho)-H_{\mathrm{min}}(\rho)=\Delta_{\mathrm{spec}}(\rho)$ (34)
@@ A.5 Combining and Rearranging the Inequality Chain
$D_{\max}^{\varepsilon}(\rho||\sigma_{R})\geq D_{\max}(\rho||\sigma_{R})-\alpha(\varepsilon)$ (35)
Bounds of this type can be derived from the standard relationships between smoothed measures and the purified distance [DattaRenner_Ref]. Second, by combining Lemma A.2 (Eq. (32)) and Lemma A.3 (Eq. (33)), we obtain the core inequality from physical cost to a smoothed information quantity. To ensure mathematical rigor, we uniformly use the form
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\,\log\Gamma_{R}$ (36)
$\quad+c_{a}\,\log\!\bigl(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}\bigr)$
$\quad+c_{b}\,\log(1/\varepsilon)+O(1)$
$\geq D_{\max}^{\varepsilon}(\rho\parallel\sigma_{R})-c_{0}$
Next, applying the aforementioned smoothing continuity bound (in the lower-bound direction) to the right-hand side and substituting the entropy identity proven in Proposition A.4 (Eq. (34)), we get:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\,\log\Gamma_{R}\geq$ $D(\rho\parallel\sigma_{R})+\Delta_{\mathrm{spec}}(\rho)$ (37)
$-\alpha(\varepsilon)-c_{0}$
$-c_{a}\,\log\!\Bigl(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}\Bigr)$
$-c_{b}\,\log(1/\varepsilon)-O(1)$
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ $\geq\frac{D(\rho\parallel\sigma_{R})}{\log\Gamma_{R}}$ (38)
$\quad-\frac{c_{1}}{\log\Gamma_{R}}\bigl[\log\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}+\log(1/\varepsilon)\bigr]$
$\quad+\frac{\Delta_{\mathrm{spec}}(\rho)}{\log\Gamma_{R}}$
@@ A.6.1 Key One-Shot Information-Theoretic Quantities
We first precisely define the core information-theoretic quantities on which the proof in this section relies. By convention: all logarithms $\log$ are base-2, errors are measured by the trace distance $(1/2)\cdot||.||_{1}$, and it is assumed by default that the support of all quantum states $\rho$ is contained within the subspace $H_{R}$ defined by the reference set RR[Tomamichel2012].
@@
    $D_{H}^{\eta}(\rho||\sigma)=-\log\min_{0\leq T\leq I,\mathrm{Tr}[T\sigma]\leq\eta}\mathrm{Tr}[(I-T)\rho]$ (39)
@@
    $D_{\max}(\rho||\sigma)=\inf\{\lambda:\rho\leq 2^{\lambda}\cdot\sigma\}$ (40)
@@ A.6.2 Core Lemmas and Relations
$D_{H}^{\eta}(E(\rho)||E(\sigma))\leq D_{H}^{\eta}(\rho||\sigma)$ (42)
$D_{\max}^{\varepsilon}(E(\rho)||E(\sigma))\leq D_{\max}^{\varepsilon}(\rho||\sigma)$ (43)
@@
$D_{H}^{\eta}(\rho||\sigma)\leq D_{\max}^{\varepsilon}(\rho||\sigma)+\kappa_{1}\cdot\log(1/\eta)+\kappa_{2}\cdot\log(1/\varepsilon)$ (44)
@@ A.6.3 Upper Bound on Smoothed Max-Relative Entropy from Circuit Complexity
$D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq L\cdot\log\Gamma_{R}+a\cdot\log(\max\{1,L\})+b\cdot\log(1/\varepsilon)$ (45)
Proof. Following the prefix-free encoding idea from Section A.2, we construct the universal semi-density operator $M=\sum_{p}2^{-|p|}\rho_{p}$. If there exists a circuit described by a program $p^{*}$ that can $\varepsilon$-approximately prepare $\rho$ (with output $\rho_{p^{*}}$), then by the definition of $M$, we have the operator inequality $M\geq 2^{-|p^{*}|}\cdot\rho_{p^{*}}$, which directly leads to $D_{\max}(\rho_{p^{*}}||M)\leq|p^{*}|$. To change the comparison object from $M$ to $\sigma_{R}$, we introduce a mixed semi-density operator $M_{\alpha}=(1-\alpha)M+\alpha\sigma_{R}$, where $\alpha\in(0,1)$. Using the scaling comparison property of max-relative entropy, we get $D_{\max}(\rho_{p^{*}}||\sigma_{R})\leq D_{\max}(\rho_{p^{*}}||M)+\log(1/(\alpha(1-\alpha)))$. Choosing $\alpha=1/2$ makes the constant term $\log(1/(\alpha(1-\alpha)))=\log(4)=2$. Combining this with the information-theoretic upper bound on $|p^{*}|$, which is $L\cdot\log\Gamma_{R}+O(\log L)$, and the continuity of smoothed entropies, we obtain Eq. (45). This proof path, via the $\sigma$-padding technique, avoids any spectral constants with explicit dependence on $d_{R}$.
@@
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ $\geq\frac{1}{\log\Gamma_{R}}\,D_{H}^{\eta}(\rho\parallel\sigma_{R})$ (46)
$\quad-\frac{c^{\prime}_{1}}{\log\Gamma_{R}}\,\biggl[\log\!\Bigl(\max\{1,C_{\mathrm{opt}}^{(\varepsilon)}(\rho)\}\Bigr)+\log(1/\varepsilon)\biggr]$
$\quad-\frac{c^{\prime}_{2}}{\log\Gamma_{R}}\,\log(1/\eta)$
Proof. Let $L=C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$. We chain together Lemma A.6.2 and Lemma A.6.3. From Eq. (44) and Eq. (45), we have: $D_{H}^{\eta}(\rho||\sigma_{R})\leq\left(L\cdot\log\Gamma_{R}+a\cdot\log(\max\{1,L\})+b\cdot\log(1/\varepsilon)\right)+\kappa_{1}\cdot\log(1/\eta)+\kappa_{2}\cdot\log(1/\varepsilon)$ By algebraically rearranging the above expression to solve for a lower bound on $L$, and defining the universal constants $c^{\prime}_{1}=\max\{a,b+\kappa_{2}\}$ and $c^{\prime}_{2}=\kappa_{1}$, we obtain Eq. (46).
@@ A.6.5 Non-Bootstrap Closed-Form Lower Bound
Eq. (46) is a bootstrap-type inequality. For practical applications, we can provide a non-recursive, closed-form lower bound for it using the method derived in Appendix C.2. Let:
$A$ $=\frac{D_{H}^{\eta}\!\bigl(\rho\,\|\,\sigma_{R}\bigr)}{\log\Gamma_{R}},$ (47)
$B$ $=\frac{1}{\log\Gamma_{R}}\!\left(c^{\prime}_{1}\log\!\frac{1}{\varepsilon}\;+\;c^{\prime}_{2}\log\!\frac{1}{\eta}\right),$ (48)
$c$ $=\frac{c^{\prime}_{1}}{\log\Gamma_{R}}$ (49)
Then Eq. (46) can be written as $L\geq A-B-c\cdot\log(\max\{1,L\})$. Its smallest solution $L^{*}$ is given by the Lambert-W function $W_{0}$:
$C_{\mathrm{opt}}^{(\varepsilon)}(\rho)$ $\geq\max\Biggl\{1,\;\frac{c}{\ln 2}\,\operatorname{W}_{0}\!\Biggl(\frac{\ln 2}{c}\,\exp\!\Bigl(\tfrac{\ln 2}{c}\,(A-B)\Bigr)\Biggr)\Biggr\}$ (50)
@@ A.6.6 Framework Consistency and Independence Check
Accounting for Errors and Compilation: The $O(\log(1/\varepsilon))$ corrections arising from smoothing and gate synthesis have been uniformly incorporated into the logarithmic correction term of Eq. (46); the constant factor from compiling parallel to serial circuits does not affect the scaling behavior and only influences the specific numerical values of the $O(1)$ coefficients $a$ and $b$.
@@ B.1 Robustness of the Theory to Support Set Leakage
$\tilde{\rho}=\frac{\Pi_{R}\cdot\rho_{\mathrm{phys}}\cdot\Pi_{R}}{\mathrm{Tr}(\Pi_{R}\cdot\rho_{\mathrm{phys}})}$ (51)
All conclusions in the main theorem, Theorem 3.1, can be directly applied to this effective state $\tilde{\rho}$. The physical meaning of this scheme is that we focus our analysis on the part of the quantum state that was successfully prepared within the target structure, treating it as the true computational goal. (Note: The above construction requires $q=\mathrm{Tr}(\Pi_{R}\rho_{\mathrm{phys}})>0$. When $q=0$ (complete leakage), the smoothed reference state scheme Sec. B.2 should be used instead.)
@@
$\tilde{\sigma}_{R}=(1-\delta)\cdot\sigma_{R}+\delta\cdot\frac{(I-\Pi_{R})}{(d-d_{R})}$ (52)
@@ B.2 Proof of the $d_{R}$-Independence of the Kolmogorov Lower Bound Constant $c_{0}$
Proposition B.3 ($d_{R}$-Independence of $c_{0}$). There exists a constant $c_{0}=O(1)$, whose magnitude depends only on the choice of the relative universal machine $U_{R}$, such that for any $\varepsilon>0$ and any $d_{R}$, it holds that $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-c_{0}$[berthiaume2001quantum].
@@
    $D_{\max}^{\varepsilon}(\rho||\sigma_{R})\leq D_{\max}^{\varepsilon}(\rho||M)+\log(1/(\alpha(1-\alpha)))$ (54)
@@
    Combining and Parameter Selection: From Eq. (53) and Eq. (54), we have:
@@
    $K_{\varepsilon}(\rho|R)\geq D_{\max}^{\varepsilon}(\rho||\sigma_{R})-2$ (55)
@@
(Note: We do not rely on any spectral comparison constants between $M$ and $\sigma_{R}$; using $\sigma$-padding (with $\alpha=1/2$) is sufficient to obtain $K_{\varepsilon}(\rho\!\mid\!R)\geq D_{\max}^{\varepsilon}(\rho\|\sigma_{R})-c_{0}$, where $c_{0}=O(1)$ and is independent of $d_{R}$, see Appendix A.3 and Section B.2.) [berthiaume2001quantum]
@@ B.3 Comparison with and Positioning Relative to Other Complexity Measures
Table 1: A multi-dimensional comparison of Reference-Contingent Complexity (RCC) with other complexity measures.
Dimension RCC Nielsen Geom. Krylov Res. Theory (Rel. Entropy) Mana / T-count
Object of measure State fn. (final state) Path fn. (dynamical) Operator growth (dynamical) Resource content of state Resource count for gate set
Structural fairness Explicit (via Ref. Set) Penalty factors dependent Partial (basis dep.) Yes (free set dep.) No (fixed model)
Relation to gate set Explicit (via $\Gamma_{R}$) Implicit (via metric) Typically independent Tied to free ops. Strong (for chosen set)
[…]
Operationality Yes (see Sec. B.3.2) Indirect (path opt.) Yes (via response fns.) Partial (free ops.) Yes (compile count)
Path dep. No Yes Yes No No
$^{\text{a}}$ LB denotes Lower Bound.